# The Role of Deep Learning Regularizations on Actors in Offline RL

### v2

@@ B.1 ReBRAC+CE
Table 3: ReBRAC+CE combined regularizations best hyperparameters.
@@ B.2 IQL
Table 5: IQL combined regularizations best hyperparameters.
@@
Appendix C Hyperparameters sensitivity
C.1 Layer Normalizations
@@ C.8 Gradient Noise
Appendix D Probabilities of Improvement for combinations of regularizations
@@
Appendix E Detailed performance comparison
@@
Table 6: Final checkpoint mean for Gym-MuJoCo tasks. SAC-DRND[Yang et al., 2024] and RORL[Yang et al., 2022] scores are taken from original works.
@@
Table 7: Final checkpoint mean for AntMaze tasks. TD3-BC-N [Beeson and Montana, 2024] scores are taken from the original work. ReBRAC+CE also stands for EF here.
@@
Table 8: Final checkpoint mean for Adroit tasks. ReBRAC+CE stands for ReBRAC+CE without hyperparameters consistency. ReBRAC+CE[Tarasov et al., 2024a] and RORL[Yang et al., 2022] scores are taken from original works.
@@
Appendix F Per-domain actor generalization
@@
Appendix G Per-domain actor internal metrics

### v3

@@ D.1 ReBRAC+CE
Table 3: ReBRAC+CE regularizations best hyperparameters.
@@ D.2 IQL
Table 5: IQL regularizations best hyperparameters.
@@
Appendix E Hyperparameters sensitivity
Sensitivity to the hyperparameters choice for all of the regularization techniques based on the experiments from subsection 4.1.
E.1 Layer Normalizations
@@ E.8 Gradient Noise
Appendix F Probabilities of Improvement for combinations of regularizations
We present rliable (Agarwal et al., 2021) Probabilities of Improvement for combined regularizations in Figure 14. In most of the cases regularizations outperform original algorithms. However, we are not able to see any clear pattern.
@@ Appendix F Probabilities of Improvement for combinations of regularizations
Appendix G Detailed performance comparison
@@
Table 6: Final checkpoint mean for Gym-MuJoCo tasks. SAC-DRND(Yang et al., 2024) and RORL(Yang et al., 2022) scores are taken from original works.
@@
Table 7: Final checkpoint mean for AntMaze tasks. TD3-BC-N (Beeson & Montana, 2024) scores are taken from the original work. ReBRAC+CE also stands for EF here.
@@
Table 8: Final checkpoint mean for Adroit tasks. ReBRAC+CE stands for ReBRAC+CE without hyperparameters consistency. ReBRAC+CE(Tarasov et al., 2024a) and RORL(Yang et al., 2022) scores are taken from original works.
@@
Appendix H Per-domain actor generalization
Here we provide experimental results from subsection 4.4 individually per D4RL domain. The only surprising observation comes from Figure 17 where the addition of regularizations reduces IQL robustness to noise perturbations.
@@ Appendix H Per-domain actor generalization
Appendix I Per-domain actor internal metrics
In this section we provide results from subsection 4.5 per D4RL domains.